# CtxMIM: Context-Enhanced Masked Image Modeling for Remote Sensing Image Understanding

### v3

@@ 4.6 Ablation Study
Figure 6: Some CtxMIM reconstructive samples on xView (Lam et al., 2018) and SpaceNet (Van Etten et al., 2018). Each row shows the ground truth, four 2-tuples (the masked image and the CtxMIM reconstruction) of different masking ratios 70%, 75%, 80%, and 85% from left to right.
5 Discussion
@@
Model scalability. SSL methods in remote sensing employ contrastive learning to pretrain foundation models and have demonstrated higher performance on different remote sensing tasks. For extensive data augmentations, these methods consider intrinsic characteristics of remote sensing images to obtain contrastive pairs, like SeCo (Manas et al., 2021), CACo (Mall et al., 2023), and SatMAE (Cong et al., 2022) leveraging the temporal information, and Geo-Aware (Ayush et al., 2021) leveraging geo-location information. However, the well-designed handcrafted data structure is a double-edged sword that limits the capabilities of foundation models to some extent. Besides, remote sensing image pairs are sensitive to significant changes in the temporal span, such as major urban development (Mall et al., 2023), natural disasters, and the external environment. Hence, CtxMIM designs a novel MIM framework by taking the natural advantage of abundant remote sensing images captured by sensors daily, making the foundation model easily extended to a larger scale in further research.

### v4

@@ IV-F Ablation Study
V Discussion
@@
Model scalability. SSL methods in remote sensing employ contrastive learning to pretrain foundation models and have demonstrated higher performance on different remote sensing tasks. For extensive data augmentations, these methods consider intrinsic characteristics of remote sensing images to obtain contrastive pairs, like SeCo [3], CACo [5], and SatMAE [39] leveraging the temporal information, and Geo-Aware [4] leveraging geo-location information. However, the well-designed handcrafted data structure is a double-edged sword that limits the capabilities of foundation models to some extent. Besides, remote sensing image pairs are sensitive to significant changes in the temporal span, such as major urban development [5], natural disasters, and the external environment. Hence, CtxMIM designs a novel MIM framework by taking the natural advantage of abundant remote sensing images captured by sensors daily, making the foundation model easily extended to a larger scale in further research.